OpenStudy (anonymous):

A rectangle has a width to length ratio of2:5. The perimeter is 322 feet. What is the area enclosed

OpenStudy (anonymous):

Let the length be 5x and breadth be 2x the perimeter is 2(5x+2x)=322 then 7x=161 x=23 the area is 5x*2x=10(x)^2 area=5290 by substituting the value of x in it!!1

OpenStudy (anonymous):

If ρ and κ are the two sides of the rectangle then ρ/κ = 2/5 (1) The perimeter is 2ρ + 2κ = 322 (2) Now time to solve.. (1) can be written as ρ = 2κ/5 and so from the second one we have 2( 2κ/5 ) + 2κ = 322 and therefore 4κ + 10κ = 322 * 5 and 14κ = 1610 so κ = 115 But remember than ρ/κ = 2/5 ? So ρ = 2κ/5 meaning ρ = 46 Finally the area is κ * p = 46 * 115 = 5290